OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

How do i find u and v?? would I use absolute value?? This is for my online class and im totally lost in this module

OpenStudy (jdoe0001):

\(\Large u\) = <-5, -4>, \(\Large v \)= <-4, -3>

OpenStudy (anonymous):

im so lost

OpenStudy (jdoe0001):

well... if you have not covered the material seems this exercise does not apply to you then

OpenStudy (anonymous):

no I have but im trying to teach myslef since its online and I cant figure it out

OpenStudy (johnweldon1993):

The "dot product" \[\large u \cdot v = |u||v|cos(\theta) \] We want to solve for the angle...so we can divide both sides by \(\large |u||v|\) \[\large cos(\theta) = \frac{u\cdot v}{|u||v|}\] And now take the inverse cos of those sides \[\large \theta = \cos^{-1}(\frac{u \cdot v}{|u||v|})\] The \(\large |u|\) and \(\large |v|\) are the "magnitudes of the vectors" \[\large |u| = \sqrt{u_x ^2 + u_y^2} \] \[\large |v| = \sqrt{v_x ^2 + v_y^2} \] Just as the dot product is calculated as \[\large u \cdot v = (u_x \times v_x + u_y \times v_y)\] That should be everything you need to do this problem! Good luck in your self-studying